OpenStudy (anonymous):

What is a40 for the arithmetic sequence presented in the table below? Hint: An = A1 + d(n - 1), where a1 is the first term and d is the common difference.

OpenStudy (anonymous):



OpenStudy (campbell_st):

ok... so you know \[A_{8} = 60 ~~~and~~~~A_{12} = 48\] so the sequence is decreasing... so the common difference is negative. so just do some simple arithmetic, remembering that the common difference is always the same value. \[\frac{(48 - 60)}{(12 - 8)} = \] and from there you can find the 1st term

OpenStudy (anonymous):

So -3 is the first term?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

no -3 is the common difference... to to find the 1st term, use this for d, and n = 8 with 60 so \[60 = a_{1} + (8 -1) \times (-3)\] now solve for your 1st term

OpenStudy (anonymous):

I got 81 for the the 1st term. Is that correct? @campbell_st

OpenStudy (campbell_st):

that's correct so now you can find the 40th term using the formula \[a_{40} = a_{1} + (40-1) \times (-3)\] hope it helps

OpenStudy (anonymous):

it did thanks!